TIP: Click on subject to list as thread! ANSI
echo: sb-nasa_news
to: All
from: Hugh S. Gregory
date: 2003-03-11 23:34:00
subject: 2\21 Pt-1 ISS On Orbit Status 21-02-2003

This Echo is READ ONLY !   NO Un-Authorized Messages Please!
 ~~~~~~~~~~~~~~~~~~~~~~~~   ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
2\21 ISS On-Orbit Status 21-02-2003
Part 1 of 2

ISS On-Orbit Status 2/21/03

All ISS systems continue to function nominally, except as noted 
previously or below. Day 90 in space for Expedition 6.

All R3 software upload activities yesterday were completed 
successfully and without hitches. The process continued today by 
transitioning primary control to the new software. [Among else, 
today's tasks were: transition primary C&C MDM to standby, followed by 
post-C&C transition reconfiguring; handover ISS attitude control to 
Russian MCS (motion control system) at about 10:40am for the software 
transition of the two GNC (guidance, navigation & control) MDMs; 
return attitude control to the U.S. segment at 1:30pm;and transition 
the INT MDMs, to conclude today's activities. Tomorrow is clean-up day 
for remaining tasks, such as installing hard drives with new software 
in the remaining PCS laptops, loading backup MDMs, etc. So far, 
everything has gone quite well.]

Due to the R3 transition activities today, the three-hour demo of EVA 
EMU operations by Bowersox and Pettit without IV (intravehicular 
crewmember) assist has been rescheduled for Monday (2/24).

The Vozdukh CO2 carbon dioxide (CO2) scrubber failed last night. Prime
suspect is BVK valve package #2, and troubleshooting is underway today 
(more tomorrow) to determine whether the failure is in the connections 
or in the valve assembly itself.

Upon loss of Vozdukh, the U.S. CDRA (CO2 removal assembly) was powered 
up and by this morning had reduced ppCO2 to 3.9 mmHg. But later today, 
CDRA failed too. It is currently operating in single-bed mode, and 
investigations by crew and ground are underway.

MCC-Moscow/TsUP has requested continuation of ppCO2 sampling by the 
crew with the U.S. CDMK (CO2 monitoring kit) until next Thursday 
(2/27), to help in calibrating the Russian measurements, now that IMV 
(intermodular ventilation) air exchange has improved. Two CDMK 
readings were taken by FE-2/SO Don Pettit today, one in the morning, 
the other in the evening.

FE-1 Nikolai Budarin continued his periodic preventive maintenance of
Russian segment (RS) atmospheric ventilation, today working in the 
DC-1 "Pirs" docking module to change out the two dust collector 
filters (PF 1 & 2).

Budarin also took photo imagery of cables on panels of the FGB "Zarya"
module. The digital images were then transferred to laptop for 
subsequent downlink. [This was in response to a previous (December 
2002) crew proposal on providing exterior protection for these cables 
in the FGB.]

Later, Budarin completed a TsUP-requested task of measuring distances
between specific mounting screws and their locations on two interior 
panels of the Service Module (SM).

Yesterday's inventory audit of EVA equipment in the Airlock by Ken 
Bowersox and Don Pettit confirmed that there is an extra SPD (spool 
positioning device) on board. This device was apparently left over 
from the 9A EVAs last October, when SPDs were installed on exterior 
ammonia QDs (quick disconnects).

Bowersox and Pettit filled out their weekly FFQ (food frequency
questionnaire), a special software log on the MEC (medical equipment
computer) to track nutritional intake.

Troubleshooting of the Foot (foot/ground reaction forces during space
flight) experiment scheduled for Ken Bowersox today has been 
postponed. Next Foot session is now planned for March 4-5. [During the 
last Foot session, six of eight data channels were not recorded. 
Ground engineers to date were unable to duplicate the failure.]

Sox performed the daily routine maintenance of the SOSh life support 
systems in the SM and the payload status checkup in the Lab 
(PCG-STES010), while Nikolai prepared the IMS database update file for 
export/downlink.

Planning is underway for a possible spacewalk by the Expedition 6 crew
during their increment. This would be some time in March.

RBVM (radiator beam valve module) temperatures are still being 
watched, but they have come down with decreasing solar Beta angle, and 
an attitude change to YVV (y-axis in velocity vector) is not required 
at this time. [As indicated by analysis (telemetry is not available), 
one thermal cover/bootie on the RBVM appears to be not correctly 
positioned. High temperatures over longer periods could potentially 
degrade the ammonia QD seals underneath.]

At 1:50pm EST, amateur radio operator Don Pettit chatted with students 
of Cowichan Secondary School of Duncan, British Columbia (Canada) 
during a 10-min. ham radio pass. [52-year old Cowichan Secondary 
School is a grade 10-12 school and one of four secondary schools 
serving the Cowichan Valley district, with a population of about 950 
students, 50 teachers and three administrators.]

The conjunction with the TRMM (Tropical Rainfall Measuring Mission, 
object #25063) predicted two days ago for tonight has "gone away" and 
will not occur.

Today's targets for the CEO (crew earth observations program) were 
Eastern Mediterranean Smog (conditions remain favorable over the basin 
of the Adriatic Sea for significant aerosol accumulations. As the crew 
tracked across central Italy they were to look either side of track 
over the water for smog plumes, especially to the left of track toward 
Venice), Eastern Great Lakes Ice and Snow (in the wake of this week's 
winter storm, looking left of track towards Lake Erie and Ontario for 
ice accumulation and the pattern of lake effect snow to their south 
and east), Western Mediterranean Smog (high pressure should still be 
holding sway over northern Italy. As ISS tracked you from the NW, the 
crew was to look left of track towards the Alps and the Po River 
valley in oblique views for smog layers), New England Snow Cover (the 
northeastern US is still digging out from the major snowstorm earlier 
this week. Looking right of track to document the extent of snow cover 
over upstate New York, Vermont, New Hampshire, and Massachusetts), and 
Chicago, Illinois (the city was probably not visible, but the crew was 
to look left of track of the eastern shore of Lake Michigan for 
wind-driven ice formations).

CEO images can be viewed at the website
http://eol.jsc.nasa.gov

(continued)

---
* Origin: SpaceBase[tm] Vancouver Canada [3 Lines] 604-473-9357 (1:153/719)
SEEN-BY: 633/267 270
@PATH: 153/719 715 7715 140/1 106/2000 633/267

SOURCE: echomail via fidonet.ozzmosis.com

Email questions or comments to sysop@ipingthereforeiam.com
All parts of this website painstakingly hand-crafted in the U.S.A.!
IPTIA BBS/MUD/Terminal/Game Server List, © 2025 IPTIA Consulting™.